# SoliDiffy: AST Differencing for Solidity Smart Contracts

### v3

@@ Abstract
Smart contracts, primarily written in Solidity, are integral to blockchain software applications, yet precise analysis and maintenance are hindered by the limitations of existing differencing tools. We introduce SoliDiffy, a novel Abstract Syntax Tree (AST) differencing tool specifically designed for Solidity. SoliDiffy enables fine-grained analysis by generating accurate and concise edit scripts of smart contracts , making it ideal for downstream tasks such as vulnerability detection, automated code repair, and code reviews. Our comprehensive evaluation on a large dataset of real-world Solidity contracts demonstrates that SoliDiffy delivers shorter and more precise edit scripts compared to state-of-the-art tools, while performing consistently in complex contract modifications. SoliDiffy is made publicly available at https://github.com/mojtaba-eshghie/SoliDiffy.
@@
AST Differencing, Solidity, Smart Contracts, Abstract Syntax Tree, Ethereum, Source Code Differencing
@@ I Introduction
Smart contracts are self-executing programs that implement real-world contracts by encoding contract terms directly into code [1, 2]. These programs are deployed on blockchain platforms like Ethereum [3], allowing for automated and trustless transactions. Solidity, a statically-typed programming language, has become the most popular choice for developing these smart contracts [4].
Developers working with Solidity can greatly benefit from a fine-grained source code differencing in several scenarios. For instance, when updating a smart contract to patch a security vulnerability, Abstract Syntax Tree (AST)-based differencing allows developers to pinpoint specific changes in the code’s structure rather than sifting through line-by-line text changes, which might miss subtle yet crucial modifications. This precision is vital when reviewing updates for correctness and security implications before deployment on an immutable blockchain. Other examples are in automated debugging [5] and program repair [6, 7, 8, 9], where tools need to detect bugs and vulnerabilities [10, 11, 12, 13, 14] and suggest changes at a syntactic level [15] ; fine-grained differencing enables these tools to generate precise edit scripts that align with the semantic intentions of the code, helping with automated and context-aware fixes. Furthermore, in code clone detection [16, 17, 18], where identifying syntactically similar but not identical code blocks is necessary, AST differencing can accurately capture variations that line-based tools overlook, thus enhancing the detection of potential code reuse or duplication issues.
Traditional text and structural differencing tools [19, 20, 21, 22, 23, 24, 25] are insufficient for smart contracts due to their inability to capture the semantic and structural details of smart contracts or their lack of support for Solidity. AST differencing for Solidity smart contracts provides a more granular approach at detection of changes at the syntactic level. To summarize, we problem we address in this paper is the fundamental limitations of existing differencing tools for Solidity developers.
In this paper, we introduce SoliDiffy, a novel AST differencing tool tailored for Solidity smart contracts. SoliDiffy contains key AST transformations specifically designed for Solidity smart contracts. SoliDiffy accurately processes and compares Solidity smart contracts, providing precise edit scripts that can be used for tasks such as vulnerability detection.
To evaluate the effectiveness of SoliDiffy, we conduct a comprehensive comparison on 354 187354187354\,187354 187 pairs of smart contracts. This dataset is founded on real-world Solidity smart contracts, including a subset of modified controlled syntactic changes (simple to complex transformation) and another subset mined from commit history of a popular smart contract repository. We assess the performance of SoliDiffy in terms of edit script accuracy. The results clearly demonstrate that SoliDiffy outperforms the only existing tool for Solidity differencing, Difftastic [21]. SoliDiffy maintains consistent effectiveness regardless of edit distance and code complexity, highlighting its suitability for advanced software engineering tasks in the blockchain domain.
To summarize, our contributions are:
  • We introduce SoliDiffy, a novel AST differencing approach specifically designed for Solidity smart contracts, addressing the limitations of existing differencing tools in accurately capturing the syntactic and semantic changes within smart contracts.
  • We design Solidity-specific AST transformation and pruning rules that enhance the precision and conciseness of edit scripts compared to a raw concrete syntax tree.
  • We conduct a comprehensive evaluation of SoliDiffy on a dataset of 354 187354187354\,187354 187 pairs of smart contracts to diff, demonstrating that it outperforms the state-of-the-art tool Difftastic in terms of edit script length and accuracy. SoliDiffy maintains consistent effectiveness regardless of edit distance and code complexity.
  • We make SoliDiffy publicly available as an open-source tool at https://github.com/mojtaba-eshghie/SoliDiffy, facilitating further research and applications in smart contract analysis.
This paper is structured as follows: Section II introduces the necessary background concepts, including ASTs and code differencing techniques. In Section III, we describe the architecture of SoliDiffy. Section IV outlines the experimental protocol including research questions that guide our investigation. Section V presents the results of our evaluation, comparing SoliDiffy with Difftastic and exploring the impact of different types and severity of syntactic changes on the performance of these tools. Section VII reviews existing work related to AST differencing and identifies gaps that SoliDiffy addresses. Finally, Section VI elaborates the lessons learnt and threats to validity of our work, and Section VIII concludes the paper.
@@
II Background
@@
II-A Syntax Trees: Abstract vs. Concrete
@@ II-B1 AST Differencing
II-B2 Edit Scripts
A common approach in AST differencing involves generating an edit script as a sequence of operations required to transform one AST into another (see Figure 3). The process typically involves two phases: generating mappings between unchanged nodes of the two ASTs and then deriving an edit script from these mappings. These edit actions—add, remove, update, and move—reflect modifications to source code. Although generating an optimal edit script is an NP-hard problem [28], this method provides a structured way to represent differences between code versions.
III SoliDiffy : AST Differencing for Solidity
@@
Figure 1 shows SoliDiffy’s architecture. SoliDiffy starts by receiving Solidity smart contracts and generating a optimized ASTs for the differencing task (Section III-A). The differencing subsystem then uses the mapping between the ASTs to perform the differencing (Section III-B).
III-A Pruning Rules
CSTs contain a wide range of unnecessary information that may pollute edit scripts. To create ASTs optimized for differencing, SoliDiffy  employs a series of transformations on the initial CSTs. This involves flattening nodes, aliasing for consistency, and pruning unnecessary elements, as follows
@@ III-A Pruning Rules
  • Ignoring: Removes extraneous nodes, such as formatting elements, that do not impact the logical structure of the code (third rule in Figure 1).
These transformation rules capture Solidity-specific constructs, and are implemented on top of pre-implemented transformation rules in Gumtree [19].
III-B Differencing Algorithm
@@
These mappings are then used to derive an edit script (Section III-C).
III-C Edit Script
SoliDiffy’s edit scripts include four standard operations: add, delete, update, and move. The differencing algorithm prioritizes producing edit scripts that are concise yet fully descriptive of the changes made.
To demonstrate edit scripts SoliDiffy generates, we use the smart contract in Figure 2. We change the identifier for a storage variable and remove a function from the source code. SoliDiffy generates the edit script in Figure 3 by identifying one update and a delete action.
1contract SimpleStorage {
2   - uint256 public num;
3   + uint256 public number_;
4 function set(uint256 _num) public {
5 num = _num;
6 }
7   - function get() public view returns (uint256) {
8   - return num;
9   - }
10}
Figure 3: Edit script generated as a result of the AST-based differencing of the code from Figure 2 using SoliDiffy. The edit actions are color-coded with delete actions presented as with red and update with orange.
III-D Implementation
SoliDiffy  uses Tree-sitter to parse Solidity source code into its syntax tree representation. Tree-sitter is an open-source parsing tool that generates CSTs for a wide range of programming languages using a modular grammar framework. We integrated the most recent version of Solidity grammar to the best of our knowledge [29]  [^1] https://github.com/JoranHonig/tree-sitter-solidity/blob/a8ed2f5d600fed77f8ed3084d1479998c649bca1/grammar.js.
IV Experimental Protocol
@@
We outline the experimental setup to answer our research questions 1–4 in Section IV-A. Figure 4 demonstrates this experimental setup, and Section IV-B elaborates on the protocol to answer the research questions.
The experiment is structured as a pipeline that begins with the selection and preparation of our seed datasets of Solidity smart contracts. The seed datasets include 1) DAppSCAN, a large dataset of smart contracts [30] and seed dataset from commit history of the Uniswap smart contracts [31] (Section IV-C). In the next phase, the contract pairs are generated from our seed datasets. Finally, we run three tools in pairs SoliDiffy , Difftastic , and Git line differencing tools according to our protocol for each research question on the generated contract pairs. The upcoming sections provide details of each stage of the experiment, including the dataset preparation, and the methodology for generating and processing source diff pairs.
IV-A Research Questions
@@
IV-B Protocol for Research Questions
IV-B1 Protocol for RQ1
To evaluate the effectiveness of SoliDiffy for AST differencing in Solidity code, we compare its performance against an existing open-source tool, Difftastic. We focus on the key metric of edit script length and success of the differencing task.
IV-B2 Protocol for RQ2
@@
IV-B3 Protocol for RQ3
The mutations used to generate the evaluation dataset range from simple syntax modifications to complex structural changes on Solidity smart contracts. We investigate the relationship between different operators and their edit distance.
IV-B4 Protocol for RQ4
@@
IV-C Datasets
IV-C1 Seed Datasets
To follow the protocols of RQ1–3, we need a dataset with a large number of Solidity source code files. For this, we use the DAppScan-source dataset [30, 32, 33]. This dataset consists of 39 9043990439\,90439 904 real-world Solidity Solidity source code files. From these 39 9043990439\,90439 904 files, $8102$ files are used for the our experiment in Figure 4, selected as follows: we remove all Solidity source files with duplicate names to ensure a more diverse dataset and reduce the potential for redundancy that could undermine the validity of our evaluation. The final dataset is available in a dedicated repository  [^2] https://github.com/SoliDiffy/SoliDiffyResults/tree/main/contracts/dataset.
For RQ4, we used commits of Uniswap v4 core smart contracts GitHub repository  [^3] https://github.com/Uniswap/v4-core.
IV-C2 Contract Pair Generation
To create contract pairs with varying levels of code alterations to the AST, one effective approach is to use a mutation testing tool [34]. Using code mutations for contract pair generation provides fine-grained code changes between versions, of better quality than what one may finds in repositories.
For this, we use mutation tool SuMo [35, 36]. It contains 44 mutation operators that are useful for evaluating various aspects of SoliDiffy [37, 38].
We used a script to invoke SuMo from the command line, generating mutants for all files in the dataset using each available mutation operator. The process involved iterating through all $44$ mutation operators, generating all possible mutations for each Solidity file, and creating up to $10$ mutated versions per file. In some cases, the actual number of generated mutants is lower than 10 due to the limited mutation opportunities in some files.
We note that some files in the dataset are incompatible with SuMo, causing crashes and preventing contract pair generation. In total, we generated 353 262353262353\,262353 262 contract pairs for differencing. The browsable version of these diff pairs is provided in our repository [^4] https://github.com/SoliDiffy/SoliDiffyResults/tree/master/contracts/mutants [^5] https://solidiffy.github.io/.
To generate contract pairs for differencing task of RQ4, we processed the commit history of Uniswap v4 core project. The contract pair generation begins by retrieving the entire commit history in chronological order using git log, followed by identifying the specific Solidity files altered in each commit through git diff-tree. For each modified file, the script extracts the version of the file at both the current and previous state and stores them. Then, a git diff between these two versions is computed using git diff, and the differences are saved to a file. This dataset of contract pairs and their differencing results are available publicly at our results repository  [^6] https://github.com/SoliDiffy/SoliDiffyResults/tree/master/uniswap-v4-diffs.
IV-D Execution Environment
@@
V Experimental Results
V-A Results for RQ1
@@
TABLE I: Effectiveness of Solidity differencing tools on our large dataset (Section IV-C1)
@@
Successful diffed pairs 339 596339596339\,596339 596 336 331336331336\,331336 331
@@
We use SoliDiffy and Difftastic to conduct a large-scale campaign of Solidity smart contract source code differencing (see Section IV). Figure 5 shows the results of running the experiment, averaged across all diff pairs of each project (with varying contract pair modification severity and different types of modifications). We present the results as a side-by-side histogram. The green bars represent SoliDiffy, while the blue bars with a hatched pattern represent Difftastic. The y axis presents the frequency of edit script length that falls into a particular bin (x axis).

### v4

@@ Abstract
Structured code differencing is the act of comparing the hierarchical structure of code via its abstract syntax tree (AST) to capture modifications. AST-based source code differencing enables tasks such as vulnerability detection and automated repair where traditional line-based differencing falls short. We introduce SoliDiffy, the first AST differencing tool for Solidity smart contracts with the ability to generate an edit script that soundly shows the structural differences between two smart-contracts using insert, delete, update, move operations. In our evaluation on 353 262353262353\,262353 262 contract pairs, SoliDiffy  achieved a $96.1\%$ diffing success rate, surpassing the state-of-the-art, and produced significantly shorter edit scripts. Additional experiments on $925$ real-world commits further confirmed its superiority compared to Git line-based differencing. SoliDiffy provides accurate representations of smart contract evolution even in the existence of multiple complex modifications to the source code. SoliDiffy is publicly available at https://github.com/mojtaba-eshghie/SoliDiffy.
@@
AST Differencing, Solidity, Smart Contracts
@@ I Introduction
Smart contracts are self-executing programs that implement real-world agreements by encoding contract terms directly into code [1, 2]. These programs operate on blockchain platforms such as Ethereum [3], enabling trustless transactions without intermediaries. Solidity, a statically-typed language, is the leading choice for smart contract development [4].
Despite the rapid adoption of smart contracts, software engineering tooling for Solidity is still scarce. In particular, the Solidity developer community lacks a robust tool for accurately tracking and analyzing code changes, essential to tasks such as automated smart-contract vulnerability detection [5, 6, 7, 8, 9] and smart-contract repair [10, 11, 12, 13, 14, 15].
While structural differencing tools exist for other programming languages [16], a good solution for smart contracts does not exist. Traditional line-based differencing tools, such as Git diff, fail to capture the hierarchical structure of smart contracts, leading to inaccurate change representations. Developers working with Solidity can benefit from fine-grained structural differencing in several scenarios. For instance, when upgrading a smart contract, AST-based differencing enables the identification of subtle structural changes that line-based methods might overlook (Section II). Similarly, automated program repair (APR) [10, 11, 12, 13, 14] relies on AST differencing to validate patches, document the transformations, and ensure semantic consistency. AST differencing also plays a crucial role in detecting code clones[17, 18, 19], where semantically similar yet syntactically different code blocks must be identified.
To address these gaps, we introduce SoliDiffy, a novel AST differencing tool tailored for Solidity smart contracts. Unlike existing line-based differencing approaches, SoliDiffy accurately captures structural modifications using Solidity-specific AST analysis. In SoliDiffy, we devise AST pruning and transformation rules to address Solidity constructs and syntax.
Next, we conduct an extensive experimental study on 354 187354187354\,187354 187 contract pairs to answer the following research questions:
  • RQ1: How does SoliDiffy compare to the state of the art? In comparison to Difftastic [20], which generates text-based edit scripts, SoliDiffy provides shorter edit scripts and successfully analyzes more contract pairs.
  • RQ2: How does SoliDiffy perform when there are multiple changes in the smart contract source code? Our analysis demonstrates that SoliDiffy remains effective and consistently generates lower edit distances regardless of the number of stacked modifications.
  • RQ3: How does the type of syntactic changes affect the performance of SoliDiffy? We find that SoliDiffy excels in handling complex structural changes, particularly large code block modifications.
  • RQ4: How does SoliDiffy perform against line differencing on real-world commit histories? Our study on Uniswap v4 [21] contracts shows that SoliDiffy provides more accurate and structured representations of code evolution compared to Git line-based differencing.
This rest if the paper is structured as follows: Section II presents a motivating example. Section III provides the necessary background. Section IV demonstrates the architecture of SoliDiffy. Section V outlines our experimental protocol, followed by results in Section VI and a discussion on key findings and threats to validity in Section VII. Section VIII reviews related work. Finally, Section IX concludes the paper.
@@
II Motivating Example
Tasks such as program repair and vulnerability detection depend on precisely understanding code modifications. Consider a scenario where a developer modifies the source code of a Solidity smart contract during an upgrade. A basic line-based differencing tool, such as Git diff, will highlight textual changes but lacks the ability to distinguish between superficial edits (e.g., renaming variables) and potentially critical changes affecting security or functionality. Such distinction becomes crucial, especially for automated tools designed to verify or reject suspicious code updates to deployed smart contracts [22, 23].
As an example, consider the SimpleStorage smart contract in Figure 1. A developer renamed the storage variable num to counter, modified its visibility from public to private, and slightly modified the formatting of existing functions. Standard line-based differencing highlights these as multiple unrelated additions and deletions, as illustrated by numerous confusing red and green lines in the figure. Such textual representation obscures the underlying structural relationship between these changes, making it challenging for reviewers or automated analysis tools to quickly grasp the meaning of the modifications.
In contrast, SoliDiffy produces an edit script (Figure 3) by precisely pinpointing the semantic modifications made in the source code. Unlike Git’s line-based diff, SoliDiffy distinguishes between changes that are semantically significant (such as variable visibility) and those that are cosmetic or formatting-related, allowing developers to more quickly understand the real intent behind source code updates.
III Background
@@
III-A Syntax Trees: Abstract vs. Concrete
@@ III-B1 AST Differencing
III-B2 Edit Scripts
A common approach in differencing is to generate an edit script as a sequence of operations required to transform one source code into another (see Figure 3). The process for AST-based edit scripts typically involves two phases: generating mappings between unchanged nodes of the two ASTs and then deriving an edit script from these mappings. These edit actions reflect modifications to source code. Although generating an optimal edit script is an NP-hard problem [26], this method provides a structured way to represent differences between code versions.
IV SoliDiffy : AST Differencing for Solidity
@@
Figure 2 shows SoliDiffy’s architecture. SoliDiffy starts by receiving a pair of Solidity smart contract source codes and generates optimized ASTs for the differencing task (Section IV-A). The differencing subsystem then uses the mapping between the ASTs to perform differencing (Section IV-B).
IV-A Pruning Rules
CSTs contain a wide range of unnecessary information that may pollute edit scripts. To create ASTs optimized for differencing, SoliDiffy  employs a series of pruning and transformations on the initial CSTs. This involves flattening nodes, aliasing for consistency, and pruning unnecessary elements, as follows:
@@ IV-A Pruning Rules
  • Ignoring: Removes extraneous nodes, such as formatting elements that do not impact the logical structure of the code (third rule in Figure 2).
We use specialized configuration of these rules to adapt to Solidity from Gumtree’s tree-sitter backend [16, 27].
IV-B Differencing Algorithm
@@
These mappings are then used to derive an edit script (Section IV-C).
1<update-node tree="visibility: public [37,43]" label="private"/>
2<update-node tree="identifier: num [44,47]" label="counter"/>
3<update-node tree="identifier: num [242,245]" label="counter"/>
4<update-node tree="identifier: num [98,101]" label="counter"/>
5<update-node tree="identifier: num [159,162]" label="counter"/>
6<update-node tree="identifier: num [292,295]" label="counter"/>
Figure 3: Edit script generated by SoliDiffy as a result of diffing task of Figure 1.
@@
IV-C Edit Script
SoliDiffy’s edit scripts include four standard operations: insert, delete, update, and move. The differencing algorithm prioritizes producing edit scripts that are concise yet fully descriptive of the changes made.
Fig. 3 shows the edit script generated by SoliDiffy for the differencing task of Fig. 1. The update action in line 1 in this edit script is the edit action to update the visibility of the state variable from public to private. The rest of the update actions propagate the renaming of num variable to counter.
IV-D Implementation
SoliDiffy uses Tree-sitter [27] to parse Solidity source code into its syntax tree representation. Tree-sitter is an open-source parsing tool that generates CSTs for a wide range of programming languages using a modular grammar framework. We integrated the most recent version of Solidity grammar [27]. [^1] https://github.com/JoranHonig/tree-sitter-solidity/blob/a8ed2f5d600fed77f8ed3084d1479998c649bca1/grammar.js
V Experimental Protocol
We outline the experimental setup to answer our research questions 1–4 in Section V-A. Figure 4 demonstrates this experimental setup, and Section V-B elaborates on the protocol to answer the research questions.
The experiment is structured as a pipeline that begins with the selection and preparation of our seed datasets of Solidity smart contracts. The seed datasets include 1) DAppSCAN, a large dataset of smart contracts [28] and seed dataset from commit history of the Uniswap smart contracts [21] (Section V-C). In the next phase, the contract pairs are generated from our seed datasets. Finally, we run three tools in pairs SoliDiffy , Difftastic , and Git line differencing tools according to our protocol for each research question on the generated contract pairs. The upcoming sections provide details of each stage of the experiment, including the dataset preparation, and the methodology for generating and processing source diff pairs.
V-A Research Questions
@@
V-B Protocol for Research Questions
V-B1 Protocol for RQ1
To evaluate how well SoliDiffy performs AST differencing for Solidity code, we compare it to an existing open-source tool, Difftastic [20]. Difftastic supports structural Solidity code differencing but does not generate fine-grained AST-based edit scripts. Difftastic generates a JSON edit script based on word-by-word replacement in the text. We compare it to the AST-based edit-scripts of SoliDiffy . In our evaluation, we focus on the key metric of edit script length (lower is better) and successful completion rate of the differencing task (higher is better).
V-B2 Protocol for RQ2
@@
V-B3 Protocol for RQ3
We generate mutations in existing Solidity contracts to evaluate the performance of SoliDiffy and Difftastic against specific types of changes. The mutations used to generate the evaluation dataset range from simple syntax modifications to complex structural changes on Solidity smart contracts. We investigate the relationship between different operators and their edit distance.
V-B4 Protocol for RQ4
@@
V-C Datasets
V-C1 Seed Datasets
To follow the protocols of RQ1–3, we need a dataset with a large number of Solidity source code files. For this, we seed synthetic modifications in the DAppScan dataset [28, 29, 30] that contains a range of real-world audited smart contract projects. This dataset consists of 39 9043990439\,90439 904 Solidity source code files. From these 39 9043990439\,90439 904 files, we select $8102$ files for the our experiment by removing all Solidity source files with duplicate names to ensure a diverse dataset and reduce the potential for redundancy that could undermine the validity of our evaluation. The final dataset is available in a dedicated repository. [^2] https://github.com/SoliDiffy/SoliDiffyResults/tree/main/contracts/dataset
For RQ4, we use the entire commit history of Uniswap v4 core smart contracts GitHub repository. [^3] https://github.com/Uniswap/v4-core
V-C2 Contract Pair Generation
To create contract pairs with varying levels of code alterations to the AST, one effective approach is to use a mutation testing tool [31]. Using mutations for difference generation provides automated, fine-grained code changes For this, we use mutation tool SuMo [32, 33], which is dedicated to Solidity. It contains 44 mutation operators that we all useSoliDiffy [34, 35].
We use a script [^4] https://github.com/mojtaba-eshghie/SoliDiffy/scripts/gen_diff_pairs.py to invoke SuMo from the command line, generating mutants for all files in the dataset using each available mutation operator. The process involves iterating through all $44$ mutation operators, generating all possible mutations for each Solidity file, and creating up to $10$ mutated versions per file. In some cases, the actual number of generated mutants is lower than 10 due to the limited mutation opportunities in some files. We note that some files in the dataset are incompatible with SuMo, causing crashes and preventing contract pair generation.
In total, we generated 353 262353262353\,262353 262 contract pairs for differencing. The browsable version of these diff pairs is provided in our repository. [^5] https://github.com/SoliDiffy/SoliDiffyResults/tree/master/contracts/mutants [^6] https://solidiffy.github.io/
To generate contract pairs for differencing task of RQ4, we processed the entire commit history of Uniswap v4 core project. The contract pair generation begins by retrieving the entire commit history of the main branch in chronological order using git log, followed by identifying the specific Solidity files altered in each commit through git diff-tree. For each modified file, the script extracts the version of the file at both the current and previous state and stores them. Then, a git diff between these two versions is computed using git diff, and the differences are saved to a file. This dataset of contract pairs and their differencing results are available publicly at our results repository. [^7] https://github.com/SoliDiffy/SoliDiffyResults/tree/master/uniswap-v4-diffs
V-D Execution Environment
@@
VI Experimental Results
VI-A Results for RQ1
@@
TABLE I: Effectiveness of Solidity differencing tools on our large dataset (Section V-C1)
@@
Successfully diffed pairs 339 596339596339\,596339 596 336 331336331336\,331336 331
@@
We use SoliDiffy and Difftastic to conduct a large-scale campaign of Solidity smart contract source code differencing (see Section V). Figure 5 shows the results of running the experiment, averaged across all diff pairs of each project (with varying contract pair modification severity and different types of modifications). We present the results as a side-by-side histogram. The green bars represent SoliDiffy, while the blue bars with a hatched pattern represent Difftastic. The y axis presents the frequency of edit script length that falls into a particular bin (x axis).